# Dynamic models of gentrification

### v1

@@ Abstract
The phenomenon of gentrification of an urban area is characterized by the displacement of lower-income residents due to rising living costs and an influx of wealthier individuals. This study presents an agent-based model that simulates urban gentrification through the relocation of three income groups – low, middle, and high – driven by living costs. The model incorporates economic and sociological theories to generate realistic neighborhood transition patterns. We introduce a temporal network-based measure to track the outflow of low-income residents and the inflow of middle- and high-income residents over time. Our experiments reveal that high-income residents trigger gentrification and that our network-based measure consistently detects gentrification patterns earlier than traditional count-based methods, potentially serving as an early detection tool in real-world scenarios. Moreover, the analysis also highlights how city density promotes gentrification. This framework offers valuable insights for understanding gentrification dynamics and informing urban planning and policy decisions.
1Institute of Information Science and Technologies (ISTI), National Research Council (CNR), Pisa, Italy
2Scuola Normale Superiore, Pisa, Italy
3Department of Computer Science, University of Pisa, Pisa, Italy
4IMT School for Advanced Studies, Lucca, Italy
5Mathematical Institute, University of Oxford, United Kingdom
@@ 1 Introduction
In this study, we present an agent-based model of gentrification, inspired by the work of Schelling on urban segregation[7]. Rather than focusing on replicating housing market dynamics, we base our analysis of gentrification on the relocation flows of citizens in a stylised urban grid. Our model is founded on the key assumption that gentrification is driven by socioeconomic inequality and additionally by differing relocation strategies across income levels. Based on insights from the literature [29, 30, 31], we assume that agents relocate according to their socioeconomic conditions[32]: low-income agents move when priced out of a neighbourhood, medium-income agents gravitate toward areas with similar economic conditions and quality of life, while high-income agents are attracted to areas undergoing economic growth where they can maximise investment returns. These contrasting behaviours, along with a heavy-tailed income distribution, are the only drivers of neighbourhood transformations. While our agent-based model builds on simple rules, it generates complex and emergent dynamics, requiring a rigorous complex systems approach to quantify and interpret the multifaceted aspects of gentrification.
Within our modeling framework, we show that gentrification emerges only when high-income residents have some mobility, even if minimal, highlighting how their movement patterns catalyse the process. We treat relocation flows of agents in our city as time-varying edges in a temporal network[33, 34], leveraging established tools from network science and human mobility research[35, 36, 37, 38, 3]. We introduce two novel measures to quantify gentrification within our theoretical framework. The first measure translates the conventional definition of gentrification into a metric based on the over-representation of middle- and high-income agents in a given area. The second measure captures the dynamics of gentrification by tracking the inflow of these agents alongside the simultaneous outflow of low-income residents. Our findings demonstrate that this dynamic measure can consistently detect gentrification earlier than traditional count-based approaches, making it a potential early-warning indicator for policymakers aiming to mitigate its impacts. Additionally, our framework can simulate the effects of various urban planning strategies and city characteristics on gentrification. Notably, we observe a direct correlation between urban density and the frequency of gentrification events. Overall, our model and measures offer a comprehensive perspective on gentrification, shedding new light on this complex urban phenomenon.
@@ 2 Agent-based model of gentrification
Modelling the city and its citizens.
The spatial distribution of the agents follows a socioeconomic radial gradient: $H$ agents predominantly occupy the city centre, $M$ agents populate the inner areas, and $L$ agents are concentrated in the periphery (Figure 1a). This mono-centric structure reflects the presence of a dominant central business district, found in many cities or metropolitan areas of varying sizes[40, 41]. Each cell (neighbourhood) $j$ has a fixed maximum capacity $K$, which limits its occupancy, that is, how many agents can stay there. When $n(j)$, the number of agents in the cell $j$, reaches maximum capacity $K$, agents are allocated to the nearest available locations. Our model operates based on two key parameters: the parameter $p_{H}$ that corresponds to the probability that a $H$ agent relocates from its current cell to a new one where the median agent-income is increasing; the parameter $\epsilon$ that is the width of the time window over which $H$ agents evaluate cell growth trends. Our model simulates a 7x7 grid city comprising 49 neighbourhoods, a scale consistent with moderately large urban areas. To ensure the robustness of our findings, we extended our analysis to an even larger urban environment (9$\times$9 grid, 81 neighbourhoods), obtaining similar results (see Supplementary Information 4).
Agent relocation dynamics.
Our model implements income-dependent relocation rules for three agent classes across a set of cells, where each cell $i$ has a maximum capacity $K$. Let $n(i)$ denote the number of agents in cell $i$ at any given time. Figure LABEL:fig:fig_0a-f illustrates the rules behind agents’ behaviours.
A $L$ agent is likely to move away from a cell if they are significantly poorer than other agents in the cell. The agent moves, with higher probability, to a cell where the income gap with current residents is smaller. The probability of a $L$ agent leaving cell $i$ at time $t$, shown in Figure LABEL:fig:fig_0a, is given by:
@@
where $\gamma^{L}_{i}(t)\in[0,1]$ represents the agent’s relative income percentile within cell $i$’s income distribution at time $t$. Upon deciding to move, $L$ agents select from the set $\mathcal{J}$ of available cells, defined as cells where $n(j)<K$. As illustrated in Figure LABEL:fig:fig_0b, the probability of selecting cell $j$ is:
@@
A $M$ agent is likely to move away when its income significantly deviates from the median income in the current cell. The agent will move to a cell where the gap with the median income is lower. Their relocation probability, visualized in Figure LABEL:fig:fig_0c, follows::
@@
where $\gamma^{M}_{i}(t)\in[0,1]$ is the $M$ agent’s relative income percentile in cell $i$. Figure LABEL:fig:fig_0d shows how their destination selection probability is determined by:
@@
A $H$ agent moves with a fixed probability $p_{H}\in[0,1]$ that reflects the frequency with which profit-orientated investments are made in the city. Let $\tilde{w}_{j}(t)$ denote the median income in cell $j$ at time $t$. $H$ agents select from set $\mathcal{H}$ of cells satisfying both $n(j)<K$ and $\phi_{j}^{\epsilon}(t)>\phi_{i}^{\epsilon}(t)$, where $\phi_{j}^{\epsilon}(t)$ is the average growth rate of median income over the past $\epsilon$ time steps, as reported in (Figure LABEL:fig:fig_0e-f):
$\phi_{j}^{\epsilon}(t)=\frac{1}{\epsilon}\sum_{\tau=t-\epsilon}^{t}[\tilde{w}_{j}(\tau)-\tilde{w}_{j}(\tau-1)].$ (5)
@@
$q^{H}_{j}(t)=\frac{\phi_{j}^{\epsilon}(t)}{\sum_{n\in\mathcal{H}}\phi_{n}^{\epsilon}(t)}.$ (6)
The simulation terminates at time $T$ when all $L$ agents can only find cells that would place them in the lowest income percentile, or after 300 time steps. We denote the termination time as T. All simulations were implemented using the Python library mesa[42]. The complete source code for model implementation and experimental procedures is available at: https://github.com/mauruscz/Gentrification
@@ 3 Measures of gentrification
Our gentrification model generates dynamic flows of agents of the three types among grid cells (see Figure LABEL:fig:fig_0g). We model these time-varying flows as a dynamic network where nodes represent grid cells and edges correspond to movements of agents of the three types. The network consists of three layers, each representing flows of $L$, $M$, or $H$ agents. This representation results in a temporal network[43, 44, 33],which is multi-layer (one per each edge type)[45] and weighted[46], defined as follows:
@@
where $V$ denotes the set of nodes (grid cells) in the network, $\alpha$ represents the layer index corresponding to $L$, $M$ or $H$ agents, and $A^{\alpha}(t)$ is the adjacency matrix corresponding to the network in layer $\alpha$ at time $t$. The matrix elements of $\mathbf{A}^{\alpha}(t)$ correspond to weighted, directed edges connecting node pairs in layer $\alpha$, representing the relocation flows of agents of type $\alpha$ between contiguous time steps of the agent-based model. For each layer $\alpha\in[L,M,H]$ at time $t$, we define three quantities: $n^{i}_{\alpha}(t)$, the number of agents in node $i$, and $s^{\text{in}}_{i,\alpha}(t)$ and $s^{\text{out}}_{i,\alpha}(t)$, the in- and out-strength of node $i$, respectively. The latter two quantify the flow of agents in $[L,M,H]$ moving to or from node $i$ between $t-1$ and $t$:
@@
Count-based measure.
@@
$\mathcal{G}_{count}^{i}(t,\Delta)$ represents the average fraction of $H$ and $M$ agents in node $i$ over the time window $[t-\Delta,t]$. We establish a significance threshold $n_{H,M}^{*}$, defined as the expected node-wise fraction of $H$ and $M$ agents under uniform random distribution across the grid. Values of $\mathcal{G}_{count}^{i}(t,\Delta)$ exceeding $n_{H,M}^{*}$ indicate an over-representation of $H$ agents. Gentrification is identified when $\mathcal{G}_{count}^{i}(t,\Delta)$ transitions from below to above this threshold. This metric functions as a node property, independent of the network edge dynamics.
@@
Network-based measure.
@@ 4 Results
We examine the impact of $H$ agent mobility on gentrification dynamics. Simulations were conducted with varying $H$ agent movement probabilities ($p_{H}$), while maintaining a fixed evaluation window of $\epsilon=20$ time steps for node growth rates.
Figure 2 illustrates the influence of $p_{H}$ on spatial gentrification patterns. At $p_{H}=0$, both $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ yield 0%, indicating complete absence of gentrification when $H$ agents are static. Introducing minimal $H$ agent mobility ($p_{H}=0.01$) triggers gentrification in 20-40% of nodes, according to both metrics. This abrupt transition highlights the critical role of $H$ agent mobility in initiating the gentrification process. As $p_{H}$ increases, we observe a monotonic rise in gentrification levels, with $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ showing similar trends but slightly different magnitudes.
Figure 2: Role of relocation rate of high-income agents $\mathbf{p_{H}}$: gentrification levels and times across 150 simulation of the model with a fixed $\epsilon=20$, $N=2^{12}$ and varying value of $p_{H}$ according to $\mathcal{G}_{bin}^{city}$ (a) and to the network based measure $\mathcal{G}_{net}^{city}$ (b).
@@
To verify the consistency with which $\mathcal{G}_{net}$ peaks precede $\mathcal{G}_{bin}$ transitions across our simulations, we conduct a lagged cross-correlation analysis between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ time series. Specifically, we first compute the cross-correlation for each node in the city grid, then average these cross-correlations across all nodes. This process is repeated for 150 independent model runs, and the results are again averaged to obtain the final cross-correlation profile, reported in Figure 3c. We compute cross-correlations for lags $\tau\in[-15,+15]$ (see Methods for details), with significance tested against a null distribution. The highest correlation at $\tau=-9$ shows that $\mathcal{G}_{net}$ peaks systematically anticipate $\mathcal{G}_{bin}$ transitions by approximately 9 time steps. We further validate these findings by developing a null model where agent movements are completely stochastic. The anticipatory behaviour of $\mathcal{G}_{net}$ relative to $\mathcal{G}_{bin}$ transitions disappears in the randomised version of the model (see Supplementary Information 2 for detailed analysis). This result support our main findings, demonstrating that $\mathcal{G}_{net}$ reliably anticipates $\mathcal{G}_{bin}$ only in scenarios where agents’ movements are influenced by their income distribution of in the city neighbourhoods.
@@
Figure 3: Evolution of a gentrification event for a representative node ($p_{H}=0.1$, $\epsilon=20$, $\Delta=15$) and $N=2^{12}$. a): Comparison of $\mathcal{G}_{net}$ (in gold) and $\mathcal{G}_{bin}$ (in black) measures over time highlighting two key moments : during and after the peak in $\mathcal{G}_{net}$, i.e., $t_{peak}$ and $t_{post}$, respectively. b): Relocation in- and out-flows during (top), and after (bottom) the peak of $\mathcal{G}_{net}$. Red arrows correspond to $L$ agent outflows, blue arrows correspond to $M$ and $H$ agent inflows within $\Delta$. (top) Peak: Simultaneous $L$ outflow and $M+H$ inflow. (bottom) Post-peak: Diminishing $L$ outflow, intensifying $M$ and $H$ inflow. Flows represent gross out- and in- strengths. c): Average cross-correlation $\langle R\rangle$ between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ across all nodes and 150 model runs, as a function of lag $\tau$. Grey area shows 95% confidence interval; dotted line indicates null distribution mean.
@@
We analyse the relationship between $\mathcal{G}_{net}$ and socioeconomic neighborhood dynamics in Figure 4. Nodes exhibiting $\mathcal{G}_{net}$ peaks are colour-coded, while those where $\mathcal{G}_{net}$ remains at zero are depicted in grey (Figure 4a). Median richness time series reveal a starting trimodal distribution: high-richness central nodes, intermediate-richness nodes, and low-richness nodes (Figure 4b). Several nodes transition from low to intermediate richness, coinciding with $\mathcal{G}_{net}$ peaks. A notable exception (pink curve) displays a sharp richness increase followed by a steep decline, ultimately transitioning from low to middle income. $\mathcal{G}_{net}$ correctly detects gentrification only after the moment this curve transitions from low to middle income, disregarding the earlier fluctuations. The overall correspondence between the colour-coded curves in both figures validates the capacity of $\mathcal{G}_{net}$ to identify gentrification solely based on relocation patterns.
@@
Figure 4: $\mathcal{G}_{net}$ peaks capture median richness transitions in neighbourhoods: a) Time series of $\mathcal{G}_{net}(t,\Delta)$, with fixed values of the time window width $\Delta=15$, of only the nodes that undergo sharp peaks of $\mathcal{G}_{net}$. b) Time series of the median richness (median of agents’ income) of all nodes in the network: the colored curves correspond to the same nodes whose time series of $\mathcal{G}_{net}$ is presented in a). The black vertical line in both plots indicates the initial time step when $H$ agents become eligible to relocate, occurring $\epsilon$ steps after the simulation’s start.
@@
In Figures 5a-c we show the relationship between urban density and gentrification levels across different values of the model parameter $p_{H}$. We conduct 150 simulations for each configuration, with fixed values of the parameters $\epsilon=20$ and $\Delta=10$. To model increasing urban density, we varied the number of agents ($N$) present in the grid, while keeping the grid dimension constant at 7x7 along with the capacity of individual nodes $K$.

### v3

@@ Abstract
The phenomenon of gentrification of an urban area is characterized by the displacement of lower-income residents due to rising living costs and an influx of wealthier individuals. This study presents an agent-based model that simulates urban gentrification through the relocation of three income groups – low, middle, and high – driven by living costs. The model incorporates economic and sociological theories to generate realistic neighborhood transition patterns. We introduce a temporal network-based measure to track the outflow of low-income residents and the inflow of middle- and high-income residents over time. Our experiments reveal that high-income residents trigger gentrification and that our network-based measure consistently detects gentrification patterns earlier than traditional count-based methods, potentially serving as an early detection tool in real-world scenarios. Moreover, the analysis highlights how city density promotes gentrification. This framework offers valuable insights for understanding gentrification dynamics and informing urban planning and policy decisions.
keywords:
Gentrification; Urban Dynamics; Temporal Networks
@@ 1 Introduction
In this study, we present an agent-based model of gentrification, inspired by the work of Schelling on urban segregation[46]. Rather than focusing on replicating housing market dynamics, we base our analysis of gentrification on the relocation flows of citizens in a stylised urban grid. Our model is founded on the key assumption that gentrification is driven by socioeconomic inequality and additionally by differing relocation strategies across income levels. As Nieuwenhuis et al.[36] found, in high-inequality contexts, low-income groups are confined to deprived areas while high-income groups maintain spatial dominance, creating a “vicious circle” that perpetuates inequality. Based on further insights from the literature[8, 3, 19], we assume that agents relocate according to their socioeconomic conditions[31]: low-income agents move when priced out of a neighbourhood, medium-income agents gravitate toward areas with similar economic conditions and quality of life, while high-income agents are attracted to areas undergoing economic growth where they can maximise investment returns. These contrasting behaviours, along with a heavy-tailed income distribution, are the only drivers of neighbourhood transformations. While our agent-based model builds on simple rules, it generates complex and emergent dynamics, requiring a rigorous complex systems approach to quantify and interpret the multifaceted aspects of gentrification.
Within our modeling framework, we show that gentrification emerges only when high-income residents have some mobility, even if minimal, highlighting how their movement patterns catalyse the process. We treat relocation flows of agents in our city as time-varying edges in a temporal network[32, 15], leveraging established tools from network science and human mobility research[4, 37, 40, 33, 41]. We introduce two novel measures to quantify gentrification within our theoretical framework. The first measure translates the conventional definition of gentrification into a metric based on the over-representation of middle- and high-income agents in a given area. The second measure captures the dynamics of gentrification by tracking the inflow of these agents alongside the simultaneous outflow of low-income residents. Our findings demonstrate that this dynamic measure can consistently detect gentrification earlier than the more intuitive, count-based approaches, making it a potential early-warning indicator for policymakers aiming to mitigate its impacts. Additionally, our framework can simulate the effects of various urban planning strategies and city characteristics on gentrification. Notably, we observe a direct correlation between urban density and the frequency of gentrification events. Overall, our model and measures offer a comprehensive perspective on gentrification, shedding new light on this complex urban phenomenon.
@@ 2 Agent-based model of gentrification
The spatial distribution of the agents follows a socioeconomic radial gradient: $H$ agents predominantly occupy the city centre, $M$ agents populate the inner areas, and $L$ agents are concentrated in the periphery (Figure 1a). This mono-centric structure reflects the presence of a dominant central business district, found in many cities or metropolitan areas of varying sizes[6, 2]. Each cell (neighbourhood) $j$ has a fixed maximum capacity $K$, which limits its occupancy, that is, how many agents can stay there. When $n(j)$, the number of agents in the cell $j$, reaches maximum capacity $K$, agents are allocated to the nearest available locations. Our model operates based on two key parameters: the parameter $p_{H}$ that corresponds to the probability that a $H$ agent relocates from its current cell to a new one where the median agent-income is increasing; the parameter $\theta$ that is the width of the time window over which $H$ agents evaluate cell growth trends. Our model simulates a 7x7 grid city comprising 49 neighbourhoods, a scale consistent with moderately large urban areas. To ensure the robustness of our findings, we extended our analysis to an even larger urban environment ($\times9$ grid, 81 neighbourhoods), as well as to a city with maximally mixed distribution of agents of the three types, where the radial gradient of income is lost, obtaining similar results (see Supplementary Information 3-5).
Figure 2: Income-profile-specific agent behaviour: The city is represented as a grid at time $t$, with three types of agents distributed across the cells. We track a low-income agent $L$ (red cell), a middle-income agent $M$ (green cell), and a high-income agent $H$ (blue cell). (a) The $L$ agent moves from its current cell $i$ with a probability $p^{L}_{i}(t)$, which is inversely proportional to its income percentile in ii. b) When $L$ moves, it relocates to a new cell $j$ with probability $q^{L}_{j}(t)$. (c) The $M$ agent moves from its current cell $i$ with probability $p^{M}_{i}(t)$, that its higher the more the income percentile of $M$ is extreme. d) When $M$ moves, it relocates to a new cell $j$ with probability $q^{M}_{j}(t)$. (e) The $H$ agent moves with a fixed probability $p_{H}$. f) The $H$ agent moves to a new cell $j$ based on its growth rate $\phi^{\theta}_{j}(t)$. At time $t+1$, the $L$ agent moves to cell $j=8$, the $M$ agent to cell $j=2$, and the $H$ agent to cell $j=6$. g) The output of our model is a temporal network, where nodes are grid cells and edges represent agent flows between cells over time. The first snapshot of the network at $t+1$ shows the flows generated by a-f). The model stops when no $L$ agent can move, resulting in no red edges in the final network snapshot at time $T$.
2.2 Agent relocation dynamics.
Our model implements income-dependent relocation rules for three agent classes across a set of cells, where each cell $i$ has a maximum capacity $K$. Let $n(i)$ denote the number of agents in cell $i$ at any given time. Figure 2a-f illustrates the rules behind agents’ behaviours.
A $L$ agent is likely to move away from a cell if they are significantly poorer than other agents in the cell. The agent moves, with higher probability, to a cell where the income gap with current residents is smaller. The probability of a $L$ agent leaving cell $i$ at time $t$, shown in Figure 2a, is given by:
@@
where $\gamma^{L}_{i}(t)\in(0,1]$ represents the agent’s relative income percentile within cell $i$’s income distribution at time $t$. Upon deciding to move, $L$ agents select from the set $\mathcal{J}$ of available cells, defined as cells where $n(j)<K$. As illustrated in Figure 2b, the probability of selecting cell $j$ is:
@@
A $M$ agent is likely to move away when its income significantly deviates from the median income in the current cell. The agent will move to a cell where the gap with the median income is lower. Their relocation probability, visualized in Figure 2c, follows::
@@
where $\gamma^{M}_{i}(t)\in[0,1]$ is the $M$ agent’s relative income percentile in cell $i$. Figure 2d shows how their destination selection probability is determined by:
@@
A $H$ agent moves from its current cell $i$ with a fixed probability $p_{H}\in[0,1]$. Let $\tilde{w}_{j}(t)$ denote the median income in cell $j$ at time $t$. $H$ agents select from set $\mathcal{H}$ of cells satisfying both $n(j)<K$ and $\phi_{j}^{\theta}(t)>\phi_{i}^{\theta}(t)$, where $\phi_{j}^{\theta}(t)$ is the average growth rate of median income over the past $\theta$ time steps, as reported in Figure 2e-f:
$\phi_{j}^{\theta}(t)=\frac{1}{\theta}\sum_{\tau=t-\theta}^{t}[\tilde{w}_{j}(\tau)-\tilde{w}_{j}(\tau-1)].$ (5)
@@
$q^{H}_{j}(t)=\frac{\phi_{j}^{\theta}(t)}{\sum_{n\in\mathcal{H}}\phi_{n}^{\theta}(t)}.$ (6)
The definition of the destination probability $q^{H}_{j}(t)$ reflects the profit-orientated nature of the movement of H agents. Cells that underwent a higher growth in the last $\Theta$ steps are thus more likely to be selected as destination. To ensure an easier interpretation of the results of our model, we impose that H agents do not operate (move) in the first $\theta$ steps, i.e., until a full evaluation of the growth-rate $\phi^{\theta}(t)$ of cells has been completed. The simulation terminates at time $T$ when all $L$ agents can only find cells that would place them in the lowest income percentile, or after 300 time steps. We denote the termination time as T. All simulations were implemented using the Python library mesa[22]. The complete source code for model implementation and experimental procedures is available at: https://github.com/mauruscz/Gentrification
@@ 3 Measures of gentrification
Our gentrification model generates dynamic flows of agents of the three types among grid cells (see Figure 2g). We model these time-varying flows as a dynamic network where nodes represent grid cells and edges correspond to movements of agents of the three types. The network consists of three layers, each representing flows of $L$, $M$, or $H$ agents. This representation results in a temporal network[18, 17, 32],which is multi-layer (one per each edge type)[23] and weighted[5], defined as follows:
@@
where $V$ denotes the set of nodes (grid cells) in the network, $\alpha$ represents the layer index corresponding to $L$, $M$ or $H$ agents, and $A^{\alpha}(t)$ is the adjacency matrix corresponding to the network in layer $\alpha$ at time $t$. The matrix elements of $\mathbf{A}^{\alpha}(t)$ correspond to weighted, directed edges connecting node pairs in layer $\alpha$, representing the relocation flows of agents of type $\alpha$ between contiguous time steps of the agent-based model. For each layer $\alpha\in[L,M,H]$ at time $t$, we define three quantities: $n^{i}_{\alpha}(t)$, the number of agents in node $i$, and $s^{\text{in}}_{i,\alpha}(t)$ and $s^{\text{out}}_{i,\alpha}(t)$, the in- and out-strength of node $i$, respectively. The latter two quantify the flow of agents in $[L,M,H]$ moving to or from node $i$ between t1t-1 and $t$:
@@
In the following subsections, we introduce two complementary metrics to capture the dynamics of gentrification. The first metric adopts a “migration stock” perspective [1], aiming to formalise the typical definition found in the social science literature, which largely emphasises changes in neighbourhood composition over time. These definitions [14, 26, 51] generally focus on the transformation of a place, such as the increasing presence of higher-income residents, rather than on the movement of people itself. To complement this, the second metric takes a “migration flow” perspective [1], shifting attention to the dynamics of displacement. It specifically measures the number of $L$ agents who leave a place as a result of being pushed out by incoming $M$ and $H$ agents.
3.1 Count-based measure.
@@
$\mathcal{G}_{count}^{i}(t,\Delta)$ represents the average fraction of $H$ and $M$ agents in node $i$ over the time window $t-\Delta,t]$. We establish a significance threshold $n_{H,M}^{*}$, defined as the expected node-wise fraction of $H$ and $M$ agents under uniform random distribution across the grid. Values of $\mathcal{G}_{count}^{i}(t,\Delta)$ exceeding $n_{H,M}^{*}$ indicate an over-representation of $H$ agents. Gentrification is identified when $\mathcal{G}_{count}^{i}(t,\Delta)$ transitions from below to above this threshold. This metric functions as a node property, independent of the network edge dynamics.
@@
3.2 Network-based measure.
@@ 4 Results
We examine the impact of $H$ agent mobility on gentrification dynamics. Simulations were conducted with varying $H$ agent movement probabilities ($p_{H}$), while maintaining a fixed evaluation window of $\theta=20$ time steps for node growth rates.
Figure 3a,b illustrates the influence of $p_{H}$ on spatial gentrification patterns. With $p_{H}=0$, both $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ yield 0%, indicating complete absence of gentrification when $H$ agents are static. Introducing minimal $H$ agent mobility ($p_{H}=0.01$) triggers gentrification in 20-40% of nodes, according to both metrics. This abrupt transition highlights the critical role of $H$ agent mobility in initiating the gentrification process. As $p_{H}$ increases, we observe a monotonic rise in gentrification levels, with $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ showing similar trends but slightly different magnitudes.
To evaluate the significance of movement rules in driving gentrification, we compare our gentrification model with two null models: (i) $R_{1}$, where agents decide whether and where to relocate with a fixed probability (50%); (ii) $R_{2}$, where only the destination choice is random (details in Supplementary Information 1). Figure 3c-g shows heatmaps of the average $\mathcal{G}_{\text{net}}^{i}(t)$ peaks per neighbourhood for the gentrification model and the two null models. Figure 3c highlights that no gentrification occurs in our model when $p_{H}=0.0$, while gentrification events persist in the null models (Figures 3d,e) even in the absence of $H$ agents’ movement. Figures 3f shows how, in our model, with $p_{H}=0.01$, gentrification events are concentrated into peripheral neighbourhoods dominated by $L$ agents. In contrast, both null models (Figures 3e,g) show widespread gentrification events, including cyclical transitions between $L$- and $M$-majority in central neighbourhoods. Supplementary Figure S1 further shows that none of the null models converge to stable configurations. The results indicate that the rules established by our model enhance the likelihood of gentrification occurring in neighbourhoods that are initially low-income. This pattern is not observed in the null models, highlighting the significance of the rules we developed for generating gentrification events, in line with the definition of gentrification found in the literature [14, 51, 48].
Figure 3: Gentrification patterns in our model: a,b) Gentrification level across 150 simulation of the model with a fixed $\theta=20$, $N=2^{12}$ and varying value of $p_{H}$ according to $\mathcal{G}_{bin}^{city}$ (a) and to the network based measure $\mathcal{G}_{net}^{city}$ (b). (c-g) Heatmaps of the average number of peaks of $\mathcal{G}_{net}^{i}(t)$ observed in any neighborhood $i$ over 150 simulations of our model (left), the $R_{2}$ null model (center) and the $R_{1}$ null model (right).
@@
To verify the consistency with which $\mathcal{G}_{net}$ peaks precede $\mathcal{G}_{bin}$ transitions across our simulations, we conduct a lagged cross-correlation analysis between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ time series. Specifically, we first compute the cross-correlation for each node in the city grid, then average these cross-correlations across all nodes. This process is repeated for 150 independent model runs, and the results are again averaged to obtain the final cross-correlation profile, reported in Figure 4c. We compute cross-correlations for lags $\tau\in[-15,+15]$ (see Methods for details), with significance tested against a null distribution. The highest correlation at $\tau=-9$ shows that $\mathcal{G}_{net}$ peaks systematically anticipate $\mathcal{G}_{bin}$ transitions by approximately 9 time steps. We further validate these findings repeating the analyses in the two randomized versions of the model. The anticipatory behaviour of $\mathcal{G}_{net}$ relative to $\mathcal{G}_{bin}$ transitions disappears in both the fully random and random-destination versions of the model (see Supplementary Figure S2). This result support our main findings, demonstrating that $\mathcal{G}_{net}$ reliably anticipates $\mathcal{G}_{bin}$ only in scenarios where agents’ movements are influenced by their income distribution of in the city neighbourhoods. More measures and parameter combination are provided in Supplementary Information 2.
@@
Figure 4: Evolution of a gentrification event for a representative node ($p_{H}=0.1$, $\theta=20$, $\Delta=15$) and $N=2^{12}$. a): Comparison of $\mathcal{G}_{net}$ (in gold) and $\mathcal{G}_{bin}$ (in black) measures over time highlighting two key moments : during and after the peak in $\mathcal{G}_{net}$, i.e., $t_{peak}$ and $t_{post}$, respectively. b): Relocation in- and out-flows during (top), and after (bottom) the peak of $\mathcal{G}_{net}$. Red arrows correspond to $L$ agent outflows, blue arrows correspond to $M$ and $H$ agent inflows within $\Delta$, node size is proportional to agents population. (top) Peak: Simultaneous $L$ outflow and $M+H$ inflow. (bottom) Post-peak: Diminishing $L$ outflow, intensifying $M$ and $H$ inflow. Flows represent gross out- and in- strengths. c): Average cross-correlation $\langle R\rangle$ between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ across all nodes and 150 model runs, as a function of lag $\tau$. Grey area shows 95% confidence interval; dotted line indicates null distribution mean.
@@
We analyse the relationship between $\mathcal{G}_{net}$ and socioeconomic neighborhood dynamics in Figure 5. Nodes exhibiting $\mathcal{G}_{net}$ peaks are colour-coded, while those where $\mathcal{G}_{net}$ remains at zero are depicted in grey (Figure 5a). We now consider the median of the income distribtion of agents populating a neighbourhood at a specific time as the richness of that neighbourhood at that time. Richness time series reveal a starting trimodal distribution: high-richness central nodes (neighbourhoods), intermediate-richness nodes, and low-richness nodes (Figure 5b). Several nodes transition from low to intermediate richness, coinciding with $\mathcal{G}_{net}$ peaks. A notable exception (pink curve) displays a sharp richness increase followed by a steep decline, ultimately transitioning from low to middle richness. $\mathcal{G}_{net}$ correctly detects gentrification only after the moment this curve transitions from low to middle richness, disregarding the earlier fluctuations. The overall correspondence between the colour-coded curves in both figures validates the capacity of $\mathcal{G}_{net}$ to identify gentrification solely based on relocation patterns.
@@
Figure 5: $\mathbf{\mathcal{G}_{net}}$ peaks capture richness transitions in neighbourhoods: a) Time series of $\mathcal{G}_{net}(t,\Delta)$, with fixed values of the time window width $\Delta=15$, of only the nodes that undergo sharp peaks of $\mathcal{G}_{net}$. b) Time series of the richness (median of agents’ income) of all nodes in the network: the colored curves correspond to the same nodes whose time series of $\mathcal{G}_{net}$ is presented in a). The black vertical line in both plots indicates the initial time step when $H$ agents become eligible to relocate, occurring $\theta$ steps after the simulation’s start.
@@
In Figures 6a-c we show the relationship between urban density and gentrification levels across different values of the model parameter $p_{H}$. We conduct 150 simulations for each configuration, with fixed values of the parameters $\theta=20$ and $\Delta=10$. To model increasing urban density, we run each simulation with a different number of agents ($N$) present in the grid, while keeping the grid dimension constant at 7x7 along with the capacity of individual nodes $K$.
@@
In addition to city-level dynamics, neighbourhood density, limited by fixed capacity, plays an interesting role. In particular, we observe that the initially less populated low-income neighbourhoods become increasingly crowded over time.This is driven by the displacement pressure exerted by higher-income agents, who push lower-income agents out of more desirable areas. As a result, low-income agents concentrate into a few affordable neighborhoods, leading to significant overcrowding. A detailed analysis of this process is provided in Supplementary Information 5.